OpenStudy (anonymous):

A circle is defined by the following equation: 3x^(2)+3y^(2)+30x+30y+150=0 What is its radius?

OpenStudy (amistre64):

divide all by 3 (x^2 +10x) + (y^2 +10y) = -50

OpenStudy (amistre64):

now complete the squares of the x and y (x^2 +10x +25) + (y^2 +10y +25) = -50-25-25 (x+5)^2 +(y+5)^2 = -100 are you sure it aint -150?

OpenStudy (amistre64):

teh radius is most likely 10

OpenStudy (amistre64):

or 0 lol

OpenStudy (anonymous):

no, in the original equation, all numbers are positive.

OpenStudy (amistre64):

-50+25+25=0 then

OpenStudy (amistre64):

looks to be 0 for some reason

OpenStudy (amistre64):

i gotta wonder, is 0 correct?

OpenStudy (anonymous):

When dividing by 3, (x^(2)+10x)+(y^(2)+10y)=-50, how do you get 10x and 10y?

OpenStudy (amistre64):

3(x^2 +10x+25)-75 +3(y^2 +10y+25)-75= -150 3(x^2 +10x+25) +3(y^2 +10y+25)= -150 +75+75 = 0

OpenStudy (amistre64):

3*10 = 30 right?

OpenStudy (anonymous):

I dont understand how you get rid of the x and y

OpenStudy (amistre64):

i dint get rid of anything :) i moved them around to work with them easier, thats all

OpenStudy (anonymous):

but if there's a variable on the left side of the equation, shouldnt there still be a variable on the right side?

OpenStudy (anonymous):

Actually, shouldn't it remain -50-25-25, and not -50+25+25?